OpenStudy (anonymous):

Find the inverse: h(x)=1/4(x-1)

OpenStudy (jamesj):

\[ h(x) = \frac{1}{4(x-1)} \] Whatever the form is, write y = h(x) and then solve for x, x = j(y). Then j will be the inverse of h.

OpenStudy (mr.math):

Your choice of the function's name is different. Never seen someone named a function j.

OpenStudy (anonymous):

Im still confused though.

OpenStudy (across):

JamesJ is telling you to solve for y\[x=\frac{1}{4(y-1)}.\]:)

OpenStudy (anonymous):

It's the reflection of \(h(x) \) across the line y=x

OpenStudy (jamesj):

Suppose f(x) = x - 1. What is the inverse that function?

OpenStudy (anonymous):

i think it would be x+1

OpenStudy (jamesj):

Yes. Now if we had written y = f(x) = x - 1, then this is the same as solving for x If y = x-1, then x = y+1. Hence the inverse function is g(x) = x+1.

OpenStudy (jamesj):

Try another. If f(x) = 2/x, what's the inverse. Write y = f(x) = 2/x. Now solve for x.

OpenStudy (anonymous):

ok let me see

OpenStudy (anonymous):

is it x=2*y or x=2/y ?

OpenStudy (jamesj):

y = 2/x xy = 2 x = 2/y.

OpenStudy (jamesj):

So what is the inverse function of f(x) = 2/x ?

OpenStudy (anonymous):

g(x)=2/y

OpenStudy (jamesj):

g(x) = 2/x. Make sure the variables match.

OpenStudy (jamesj):

Now let's verify this. If g is the inverse function, then f(g(x)) = x and g(f(x)) = x. Let's check. f(g(x)) = f(2/x) = 2/(2/x) = 2x/2 = x. So that works. Likewise g(f(x)) = g(2/x) = 2/(2/x) = 2x/2 = x.

OpenStudy (jamesj):

Now try this one. f(x) = 1/(x+1). What's the inverse of this function?

OpenStudy (anonymous):

ok let me see

OpenStudy (anonymous):

x=0

OpenStudy (jamesj):

that can't possibly be right. Try again. Follow the method Step 1: Write y = f(x) Step 2: Solve this equation for x, i.e., find the function g so that x = g(y) Step 3: The inverse function of f is now g(x) Step 4: (optional) Verify the result by checking to make sure f(g(x)) = x and g(f(x)) = x.

OpenStudy (anonymous):

x=1/(y-1) ?

OpenStudy (jamesj):

Let's check. If g(x) = 1/(x-1) is the inverse of f(x) = 1/(x+1), then f(g(x)) = x. Well f(g(x)) = f(1/(x-1)) \[ = \frac{1}{1/(x-1) - 1} \] \[ = \frac{x-1}{1 - (x-1)} \] \[ = \frac{x-1}{-x} \] well, that's not equal to x. So look like g(x) = 1/(x-1) isn't the inverse.

OpenStudy (jamesj):

Let's try that again. y = f(x) = 1/(x+1)

OpenStudy (anonymous):

my brain feels blocked

OpenStudy (jamesj):

Solve for x. y = 1/(x+1) 1/y = x + 1 x + 1 = 1/y x = 1/y - 1

OpenStudy (jamesj):

Hence the inverse function to f(x) = 1/(x+1) is g(x) = (1/x) - 1.

OpenStudy (jamesj):

Following?

OpenStudy (anonymous):

yes

OpenStudy (jamesj):

Let's verify: \[ f(g(x)) = f((1/x) - 1) = \frac{1}{(1/x - 1) + 1} = \frac{1}{1/x} = x \] So it's looking good.

OpenStudy (anonymous):

oh ! i see. you are switching the variables and sign, so its like the opposite

OpenStudy (jamesj):

Ok. Now do your problem. h(x)=1/4(x-1). Follow the steps.

OpenStudy (jamesj):

If y = h(x) y = 1/4(x-1) Now solve for x.

OpenStudy (jamesj):

\[ y = \frac{1}{4(x-1)} \]

OpenStudy (anonymous):

umm , x=1/4y+1

OpenStudy (anonymous):

im never going to understand this.

OpenStudy (jamesj):

Yes, x = (1/4y) + 1. Hence the inverse function g(x) = what?

OpenStudy (anonymous):

Inverse function g(x) = (1/4x)+1

OpenStudy (jamesj):

Correct. Now let's verify. h(x)=1/4(x-1) h(g(x)) = h((1/4x) + 1) \[ = \frac{1}{4((1/4x) + 1) - 4} = \frac{1}{1/x + 4 - 4} = \frac{1}{1/x} = x \]

OpenStudy (jamesj):

So yes, h(g(x)) = x, the identity function. Therefore g is indeed the inverse of h.

OpenStudy (anonymous):

ok great, i think i got it now

OpenStudy (jamesj):

ok

OpenStudy (anonymous):

thanks for your help